# Determinism and Asymmetry in General Relativity

### v1

@@
Abstract
@@
This paper concerns the question of which collections of general relativistic spacetimes are deterministic relative to which definitions. We begin by considering a series of three definitions of increasing strength due to Belot (1995). The strongest of these definitions is particularly interesting for spacetime theories because it involves an asymmetry condition called “rigidity” that has been studied previously in a different context (Geroch 1969; Halvorson and Manchak 2022; Dewar 2024). We go on to explore other (stronger) asymmetry conditions that give rise to other (stronger) forms of determinism. We introduce a number of definitions of this type and clarify the relationships between them and the three considered by Belot. We go on to show that there are collections of general relativistic spacetimes that satisfy much stronger forms of determinism than previously known. We also highlight a number of open questions.
1 Introduction
@@
This paper concerns the question of which collections of general relativistic spacetimes are deterministic relative to which definitions. Consider an intuitive but imprecise idea due to Montague (1974), Lewis (1983), and Earman (1986): a collection of possible worlds is “MLE deterministic” if, for any worlds in the collection that agree on some initial segment, the worlds agree entirely. Belot (1995) considers three natural ways of making MLE determinism precise. Within the discussions of the hole argument (Earman and Norton 1987), the first two definitions have long played a central role (Butterfield 1989; Melia 1999; Pooley 2021). Belot’s third definition is much stronger than the other two and has only recently become a focus of attention (Halvorson and Manchak 2022; Landsman 2023; Cudek 2023). For each definition, Belot provides a highly symmetric example that seems to be (i) deterministic according to the particular formulation but (ii) indeterministic by the lights of certain haecceitist positions. This leads him to conclude that “determinism cannot be a formal property of theories” (Belot 1995, p. 88).
@@
In a recent paper, Halvorson et al. (2025) push back on this position. They hold that Belot’s third definition is actually a promising precisification of MLE determinism that was unfortunately ignored for thirty years. They also argue that Belot’s third example is ambiguous; until it is made precise, it is neither deterministic nor indeterministic according to his third definition. It is emphasized that whether or not a theory is deterministic depends crucially on the precise formulation of the theory and the precise formulation of determinism.
@@
Here in this paper, we follow this line but focus our attention on (variants of) general relativity. Belot’s third definition is particularly interesting for spacetime theories because it involves an asymmetry condition called “rigidity” that has been studied previously in a different context (Geroch 1969; Halvorson and Manchak 2022; Dewar 2024). We explore other (stronger) asymmetry conditions that give rise to other (stronger) forms of MLE determinism. We introduce a number of definitions of this type and clarify the relationships between them and the three considered by Belot. We go on to show that there are collections of general relativistic spacetimes that satisfy much stronger forms of determinism than previously known. We also highlight a number of open questions.
2 De Dicto and De Re Determinism
@@
Let $\mathscr{U}$ be the collection of all general relativistic spacetimes $(M,g)$. Here $M$ is a smooth, connected, Hausdorff, $n$-dimensional (for $n\geq 2$) manifold and $g$ is a smooth Lorentzian metric on $M$. In order to make sense of the notion of “initial segment” in the MLE definition of determinism, we first restrict attention to the collection $\mathscr{H}\subset\mathscr{U}$ of globally hyperbolic spacetimes with a chosen temporal orientation (see Wald 1984, Section 8.3). Any $n$-dimensional globally hyperbolic spacetime $(M,g)$ admits an $(n-1)$-dimensional Cauchy surface $\Sigma\subset M$ representing all of space at a given time. One can show that $M$ can be foliated by a collection of Cauchy surfaces such that the topology of $M$ is $\Sigma\times\mathbb{R}$. For any spacetime $(M,g)\in\mathscr{H}$, we define an initial segment $U\subset M$ to be the timelike past $I^{-}[\Sigma]$ of any Cauchy surface $\Sigma$ in $(M,g)$. One can verify that any initial segment $U$ of any globally hyperbolic spacetime $(M,g)$ counts as a connected, open subset of $M$. So $(U,g)$ is a spacetime in its own right. Let us now consider two forms of MLE determinism adapted from Belot (1995).
@@
Definition 1. A collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto deterministic if, for any spacetimes $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$ and any initial segments $U\subset M$ and $U^{\prime}\subset M^{\prime}$, if there is an isometry $\varphi:U\rightarrow U^{\prime}$, then there is an isometry $\psi:M\rightarrow M^{\prime}$.
@@
Definition 2. A collection $\mathscr{C}\subseteq\mathscr{H}$ is de re deterministic if, for any spacetimes $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$ and any initial segments $U\subset M$ and $U^{\prime}\subset M^{\prime}$, if there is an isometry $\varphi:U\rightarrow U^{\prime}$, then there is an isometry $\psi:M\rightarrow M^{\prime}$ such that $\psi_{|U}=\varphi$.
@@
These two forms of determinism (or close variations) go by different names and have long played a central role in discussions of the hole argument. The first is called “Dm2” in (Butterfield 1989), “Definition 1” in (Belot 1995), “Lewis’s analysis” in (Melia 1999), “Det2” in (Pooley 2021), and “D1” in (Halvorson et al. 2025). The second is called “Definition 2” in (Belot 1995), the “second resolution” in (Melia 1999), “Dm2+” in (Cudek 2023), and “D2” in (Halvorson et al. 2025). Here, we use the names “de dicto” and “de re” from (Dewar 2016, 2024). The two forms of determinism capture slightly different senses in which a collection of spacetimes might be considered deterministic. Clearly, de re determinism implies de dicto determinism. They differ in how one understands “agreement of worlds” in the MLE formulation of determinism. To better understand the definitions, consider the following example.
@@
Example 1. Consider two-dimensional Minkowski spacetime in standard $(t,x)$ coordinates and let $(M,g)$ be the $t<0$ region. The collection $\mathscr{C}=\{(M,g)\}$ is de dicto deterministic. This follows easily since there is only one spacetime in the collection. But $\mathscr{C}$ is not de re deterministic. To see this, consider the initial segments $U,U^{\prime}\subset M$ given by the $t<-2$ and $t<-1$ regions respectively. Clearly there is a time translation isometry $\varphi:U\rightarrow U^{\prime}$ defined by $\varphi(t,x)=(t+1,x)$. But there can be no corresponding global isometry $\psi:M\rightarrow M$ such that $\psi_{|U}=\varphi$ because of the $t=0$ “edge” of the spacetime (see Figure 1).
@@
Figure 1: Two copies of the spacetime $(M,g)$ from Example 1. The time translation isometry $\varphi$ from the initial segment $U$ to the initial segment $U^{\prime}$ does not extend to a global isometry because of the $t=0$ “edge” of the spacetime (dotted line).
Example 1 illustrates a curious state of affairs. The agreement of the initial segments $U$ and $U^{\prime}$ given by $\varphi$ does not extend to a global agreement. So one form of MLE determinism (de re) does not hold. And yet there is global agreement given by the identity map on $M$ showing that another form of MLE determinism (de dicto) does hold. A similar situation arises in the “collapsing columns” example due to Wilson (1993) often used to distinguish the two definitions. Here is Belot (1995, p. 191):
“But now consider a world $W$, in which centrally loaded columns collapse by buckling, which contains nothing but: (i) a large, homogeneous, perfectly spherical planet; (ii) a relatively small perfectly cylindrical column resting on the planet so that its axis of symmetry is normal to the planet at the point of contact; (iii) and a cone, which is moving through space in such a way that at time $t=0$ its apex will impact the column at the exact center of its top surface, with sufficient force to cause the column to buckle. We also assume that W has a Newtonian spacetime structure; that the laws concerning motion are completely deterministic; and that the law governing column collapse fully determines the shape that the column assumes upon collapse.”
The classical analogue to de dicto determinism is satisfied by the singleton collection consisting of the collapsing columns world $W$. However, by the lights of certain haecceitist positions, there is indeterminism present in the example: because of the symmetries of the $t<0$ initial segment, the columns “could have toppled in a different direction” (Melia 1999, p. 649). Accordingly, the de dicto form of MLE determinism is considered not strong enough. With respect to the classical analogue of de re definition however, the collapsing columns example fails to be de re deterministic. In this way, the example is similar to Example 1. Stepping back, we have the following.
@@
Proposition 1. If a collection $\mathscr{C}\subseteq\mathscr{H}$ is de re deterministic, then it is de dicto deterministic; the other direction does not hold.
@@
Which collections of $\mathscr{C}\subseteq\mathscr{H}$ satisfy de re determinism? Consider the collection $\mathscr{V}^{+}\subset\mathscr{H}$ of four-dimensional, inextendible, globally hyperbolic, vacuum solutions to Einstein’s equation. Let $(M,g)$ and $(M^{\prime},g^{\prime})$ be any spacetimes in $\mathscr{V}^{+}$ and let $U\subset M$ and $U^{\prime}\subset M^{\prime}$ be any initial segments. If there is an isometry $\varphi:U\rightarrow U^{\prime}$, it follows from a theorem by Choquet-Bruhat and Geroch (1969) that there is an isometry $\psi:M\rightarrow M^{\prime}$ such that $\psi_{|U}=\varphi$. Thus, we have the following which captures a sense in which large swaths of general relativistic spacetimes of physical interest are deterministic.
Proposition 2. Any collection $\mathscr{C}\subseteq\mathscr{V}^{+}$ is de re (and thus de dicto) deterministic.
3 De Re + Uniqueness
@@
Given a collection of $\mathscr{C}\subseteq\mathscr{H}$, both the de dicto and de re forms of MLE determinism require that for any spacetimes $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$ and any initial segments $U\subset M$ and $U^{\prime}\subset M^{\prime}$, if there is an isometry $\varphi:U\rightarrow U^{\prime}$, then a certain type of isometry $\psi:M\rightarrow M^{\prime}$ must exist. In the natural way, one can strengthen each definition by adding a uniqueness clause to $\psi$. Indeed, the third form of MLE determinism considered by Belot (1995) is de re determinism with uniqueness: worlds must agree in only one way.
@@
Curiously, this definition has received relatively little attention in the hole argument literature until just recently. De re* determinism is called “Definition 3” in (Belot 1995) and “Dm2++” in (Cudek 2023). It is endorsed by Halvorson et al. (2025) who call it “D3” determinism. The definition is also briefly considered by Halvorson and Manchak (2022) and Landsman (2023). Why might one be interested in de re* determinism? It has been argued that de re determinism is not strong enough to count all instances of indeterminism as such. Belot (1995) introduces his own symmetric example which leads to the formulation of his Definition 3 (de re* determinism). We will return to a similar example in due course. For now let us turn to another case: classical Leibnizian spacetime (Earman 1989). One can verify that the spacetime (considered as a singleton collection) counts as deterministic under the classical analogue of the de re formulation. And yet there are reasons to think that such a spacetime ought to be considered indeterministic. Here is Earman (1977, p. 96):
@@
“If the history of a particle is represented by a timelike world line (i.e. a world line which is everywhere oblique to the planes of simultaneity) on [Leibnizian spacetime], then determinism cannot hold. For among the automorphisms [of Leibnizian space-time] are those which are the identity on the portion of [Leibnizian spacetime] on or below some given time slice but which differ from the identity above; such a mapping leaves fixed the entire past history of the particles while changing their future behavior. Since the automorphisms of the space-time should be symmetries of the dynamical laws (whatever they are), there will be two solutions which describe the same particle histories for all past times but which describe different future behaviors.”
@@
Whether or not Leibnizian spacetime (or theories of matter set therein) ought to count as deterministic in some sense is an interesting question (see Stein 1977; Weatherall 2020). We do not enter the debate here. At this stage, we emphasize that we are not interested in capturing, once and for all, a single notion of “determinism” within within spacetime theories. There are many interesting forms of determinism to explore (Doboszewski 2019; Smeenk and Wüthrich 2021). In a pluralistic way, we are simply interested in mapping out which collections of general relativistic spacetimes are deterministic relative to which definitions. Here, we draw attention to the fact that there are definitions of determinism – in particular de re* – which are strong enough to capture Earman’s sense in which Leibnizian spacetime is indeterministic.
@@
Let us now return to general relativity. The isometry uniqueness clause which is added to de re determinism to obtain de re* determinism has the effect of requiring a certain type of spacetime asymmetry known as “rigidity” (Geroch 1969) which has recently become a focus of attention in the hole argument literature (Halvorson and Manchak 2022; Menon and Read 2023; Dewar 2024). Consider the following definition.
@@
Definition 4. A collection $\mathscr{C}\subseteq\mathscr{U}$ is rigid if, for any $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$ and any open set $O\subset M$, if there are isometries $\varphi,\psi:M\rightarrow M^{\prime}$ such that $\varphi_{|O}=\psi_{|O}$, then $\varphi=\psi$.
@@
If a collection of spacetimes is rigid, then for any pair of spacetimes in the collection, if there are isometries from one to the other that agree on an open set, they must agree entirely. It is easily verified that any collection $\mathscr{C}\subseteq\mathscr{H}$ that is both rigid and de re deterministic must be de re* deterministic. Indeed, one can think of the rigidity condition as a type of “dual” to de re determinism: the latter requires the existence of a certain isometry $\psi:M\rightarrow M^{\prime}$ while the former requires (because initial segments are open sets) the uniqueness of such an isometry (Dewar 2024). One can use a general rigidity result due to Geroch (1969, Theorem A1) to show the following (Halvorson and Manchak 2022, Theorem 1).
Proposition 3. Any collection $\mathscr{C}\subseteq\mathscr{U}$ is rigid.
@@
One way to understand the significance of the proposition is as follows: within the context of general relativity, there can be no collection of spacetimes that is de re deterministic but not de re* deterministic (cf. Cudek 2023, footnote 46). To distinguish between these two notions of determinism, one has to look to certain classical spacetime theories (eg. the Leibnizian case mentioned above) or non-standard general relativity such as its non-Hausdorff generalization (Manchak and Barrett 2023). From Proposition 3, we have the following.
Corollary 1. A collection $\mathscr{C}\subseteq\mathscr{H}$ is de re deterministic if and only if it is de re* deterministic.
@@
Recall that $\mathscr{V}^{+}\subset\mathscr{H}$ is the collection of four-dimensional, inextendible, globally hyperbolic, vacuum solutions of Einstein’s equation. From Corollary 1 and Proposition 2, we have the following result which shows that large swaths of general relativistic spacetimes of physical interest satisfy a much stronger form of determinism than previously known (cf. Landsman 2022, footnote 23).
Corollary 2. Any collection $\mathscr{C}\subseteq\mathscr{V}^{+}$ is de re* deterministic.
4 De Dicto + Uniqueness

### v2

@@
Determinism and Asymmetry in General Relativity
JB Manchak, TW Barrett, HP Halvorson, JO Weatherall
@@
This paper concerns the question of which collections of general relativistic spacetimes are deterministic relative to which definitions. We begin by considering a series of three definitions of increasing strength due to Belot. The strongest of these definitions is particularly interesting for spacetime theories because it involves an asymmetry condition called ‘rigidity’ that has been studied previously in a different context by Geroch. We show how Belot’s strongest definition is connected to a formal result emphasized by Weatherall in his paper on the hole argument. We then go on to explore other (stronger) asymmetry conditions that give rise to other (stronger) forms of determinism. We introduce a number of definitions of this type and clarify the relationships between them and the three considered by Belot. We show that there are collections of general relativistic spacetimes that satisfy much stronger forms of determinism than previously known. We also highlight a number of open questions.
1. Introduction
@@
This paper concerns the question of which collections of general relativistic spacetimes are deterministic relative to which definitions. Consider an intuitive idea developed in various ways by Montague ([1974), Lewis ([1983]), and Earman ([1986]): a collection of possible worlds is ‘MLE deterministic’ if, for any worlds in the collection that agree on some initial segment, the worlds agree entirely. Belot ([1995]) considers three natural ways of making MLE determinism precise. Within the discussions of the hole argument (Earman and Norton [1987]), the first two definitions have long played a central role (Butterfield [1989]; Melia [1999]; Pooley [2021]). Belot’s third definition is much stronger than the other two and has only recently become a focus of attention (Landsman [2023]; Cudek [2023], Halvorson and Manchak [2025]). For each definition, Belot provides a highly symmetric example that seems to be (i) deterministic according to the particular formulation but (ii) indeterministic by the lights of certain haecceitist positions. This leads him to conclude that all three definitions are inadequate.
@@
In a recent paper, Halvorson et al. ([unpublished]) push back on this position. They hold that Belot’s third definition is actually a promising precisification of MLE determinism that was unfortunately ignored for thirty years. They also argue that Belot’s third example is ambiguous; until it is made precise, it is neither deterministic nor indeterministic according to his third definition. It is emphasized that whether or not a theory is deterministic depends crucially on the precise formulation of the theory and the precise formulation of determinism.
@@
Here in this paper, we follow this line but focus our attention on (variants of) general relativity. Belot’s third definition is particularly interesting for spacetime theories because it involves an asymmetry condition called ‘rigidity’ that has been studied previously in a different context (Geroch [1969]; Halvorson and Manchak [2025]; Dewar [2025]). We show how this definition is connected to a formalresult emphasized by Weatherall ([2018]) in his paper on the hole argument. We then explore other (stronger) asymmetry conditions that give rise to other (stronger) forms of MLE determinism. We introduce a number of definitions of this type and clarify the relationships between them and the three considered by Belot. We go on to show that there are collections of general relativistic spacetimes that satisfy much stronger forms of determinism than previously known. We also highlight a number of open questions.
2. De Dicto and De Re Determinism
@@
Let $\mathscr{U}$ be the collection of all general relativistic spacetimes $(M,g)$. Here $M$ is a smooth, connected, Hausdorff, $n$-dimensional (for $n\geq 2$) manifold and $g$ is a smooth Lorentzian metric on $M$. In order to make sense of the notion of ‘initial segment’ in the MLE definition of determinism, we first restrict attention to the collection $\mathscr{H}\subset\mathscr{U}$ of globally hyperbolic spacetimes with a chosen temporal orientation (see Wald [1984]). Any $n$-dimensional globally hyperbolic spacetime $(M,g)$ admits an (n1)(n-1)-dimensional Cauchy surface $\Sigma\subset M$ representing all of space at a given time. One can show that $M$ can be foliated by a collection of Cauchy surfaces such that the topology of $M$ is $\Sigma\times\mathbb{R}$. For any spacetime $(M,g)\in\mathscr{H}$, we define an initial segment $U\subset M$ to be the timelike past $I^{-}[\Sigma]$ of any Cauchy surface $\Sigma$ in $(M,g)$. One can verify that any initial segment $U$ of any globally hyperbolic spacetime $(M,g)$ counts as a connected, open subset of $M$. So $(U,g)$ is a spacetime in its own right. Let us now consider two forms of MLE determinism adapted from Belot ([1995]):
@@
Definition 1: A collection $\mathscr{C}\subseteq\mathscr{H}$ is de dicto deterministic if, for any spacetimes $(M,g)$, $(M^{\prime},g^{\prime})\in\mathscr{C}$ and any initial segments $U\subset M$ and $U^{\prime}\subset M^{\prime}$, if there is an isometry $\varphi:U\rightarrow U^{\prime}$, then there is an isometry $\psi:M\rightarrow M^{\prime}$.
@@
Definition 2: A collection $\mathscr{C}\subseteq\mathscr{H}$ is de re deterministic if, for any spacetimes $(M,g)$, $(M^{\prime},g^{\prime})\in\mathscr{C}$ and any initial segments $U\subset M$ and $U^{\prime}\subset M^{\prime}$, if there is an isometry $\varphi:U\rightarrow U^{\prime}$, then there is an isometry $\psi:M\rightarrow M^{\prime}$ such that $\psi_{|U}=\varphi$.
@@
These two forms of determinism (or close variations) go by different names and have long played a central role in discussions of the hole argument. The first is called ‘Dm2’ in (Butterfield [1989]), ‘Definition 1’ in (Belot [1995]), ‘Lewis’s analysis’ in (Melia [1999]), ‘Det2’ in (Pooley [2021]), and ‘D1’ in (Halvorson et al. [unpublished]). The second is called ‘Definition 2’ in (Belot [1995]), the ‘second resolution’ in (Melia [1999]), ‘Dm2+’ in (Cudek [2023]), and ‘D2’ in (Halvorson et al. [unpublished]). Here, we use the names ‘de dicto’ and ‘de re’ from (Dewar [2016], [2025]). The two forms of determinism capture slightly different senses in which a collection of spacetimes might be considered deterministic. Clearly, de re determinism implies de dicto determinism. They differ in how one captures ‘agreement of worlds’ in the MLE formulation of determinism. To better understand the definitions, consider the following example.
@@
Example 1: Consider two-dimensional Minkowski spacetime in standard $(t,x)$ coordinates and let $(M,g)$ be the $t<0$ region (cf. Roberts [2020]; Manchak [2020]). The collection $\mathscr{C}=\{(M,g)\}$ is de dicto deterministic. This follows easily since there is only one spacetime in the collection. But $\mathscr{C}$ is not de re deterministic. To see this, consider the initial segments $U,U^{\prime}\subset M$ given by the t<2t<-2 and t<1t<-1 regions respectively. Clearly there is a time translation isometry $\varphi:U\rightarrow U^{\prime}$ defined by $\varphi(t,x)=(t+1,x)$. But there can be no corresponding global isometry $\psi:M\rightarrow M$ such that $\psi_{|U}=\varphi$ because of the $t=0$ ‘edge’ of the spacetime (see Figure 1).
@@
Figure 1: Two copies of the spacetime $(M,g)$ from Example 1. The time translation isometry $\varphi$ from the initial segment $U$ to the initial segment $U^{\prime}$ does not extend to a global isometry because of the $t=0$ ‘edge’ of the spacetime (dotted line).
From Example 1, we have the following:
@@
Proposition 1: If a collection $\mathscr{C}\subseteq\mathscr{H}$ is de re deterministic, then it is de dicto deterministic; the other direction does not hold.
Example 1 illustrates a curious state of affairs. The agreement of the initial segments $U$ and $U^{\prime}$ given by $\varphi$ does not extend to a global agreement. So one form of MLE determinism (de re) does not hold. And yet there is global agreement given by the identity map on $M$ showing that another form of MLE determinism (de dicto) does hold. A similar situation arises in the ‘collapsing columns’ example due to Wilson ([1993]) often used to distinguish the two definitions. Here is Belot ([1995], p. 191):
But now consider a world $W$, in which centrally loaded columns collapse by buckling, which contains nothing but: (i) a large, homogeneous, perfectly spherical planet; (ii) a relatively small perfectly cylindrical column resting on the planet so that its axis of symmetry is normal to the planet at the point of contact; (iii) and a cone, which is moving through space in such a way that at time $t=0$ its apex will impact the column at the exact center of its top surface, with sufficient force to cause the column to buckle. We also assume that W has a Newtonian spacetime structure; that the laws concerning motion are completely deterministic; and that the law governing column collapse fully determines the shape that the column assumes upon collapse.
The classical (i.e. non-relativistic) analogue to de dicto determinism is satisfied by the singleton collection consisting of the collapsing columns world $W$. However, by the lights of certain haecceitist positions, there is indeterminism present in the example: because of the symmetries of the $t<0$ initial segment, the columns ‘could have toppled in a different direction’ (Melia [1999], p. 649). Accordingly, the de dicto form of MLE determinism is considered not strong enough. With respect to the classical analogue of the de re definition however, the collapsing columns example fails to be de re deterministic. To see this, note that because of the symmetries of the $t<0$ portion of $W$, there will be a rotational isomorphism of the model (that preserves both the metric and matter content) from this initial segment to itself that takes any spatial direction into any other. But unless this isomorphism is the identity map, it cannot be extended to all of $W$ since the direction of collapsed column after $t=0$ must be taken into itself. We note that this example can be considered a violation of Curie’s principle which states that ‘when certain causes produce certain effects, the elements of symmetry of the causes must be found in the produced effects’ (Curie [1894], p. 401). For more on Curie’s principle and general relativity, see Earman ([2007]), Roberts ([2016]), and Manchak ([unpublished]).
@@
Which collections of $\mathscr{C}\subseteq\mathscr{H}$ satisfy de re determinism? Consider the collection $\mathscr{V}^{+}\subset\mathscr{H}$ of four-dimensional, inextendible, globally hyperbolic, vacuum solutions to Einstein’s equation. Let $(M,g)$ and $(M^{\prime},g^{\prime})$ be any spacetimes in $\mathscr{V}^{+}$ and let $U\subset M$ and $U^{\prime}\subset M^{\prime}$ be any initial segments. If there is an isometry $\varphi:U\rightarrow U^{\prime}$, it follows from a theorem by Choquet-Bruhat and Geroch ([1969]) that there is an isometry $\psi:M\rightarrow M^{\prime}$ such that $\psi_{|U}=\varphi$. Thus, we have the following which captures a sense in which significant swathes of general relativistic spacetimes of physical interest are deterministic. (We note here that the result can be generalized to include collections of non-vacuum spacetimes as well. The interested reader can consult Wald ([1984], pp. 266-267) for details.)
Proposition 2: Any collection $\mathscr{C}\subseteq\mathscr{V}^{+}$ is de re (and thus de dicto) deterministic.
3. De Re + Uniqueness
@@
Given a collection of $\mathscr{C}\subseteq\mathscr{H}$, both the de dicto and de re forms of MLE determinism require that for any spacetimes $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$ and any initial segments $U\subset M$ and $U^{\prime}\subset M^{\prime}$, if there is an isometry $\varphi:U\rightarrow U^{\prime}$, then a certain type of isometry $\psi:M\rightarrow M^{\prime}$ must exist. In the natural way, one can strengthen each definition by adding a uniqueness clause to $\psi$. Indeed, the third form of MLE determinism considered by Belot ([1995]) is de re determinism with uniqueness: worlds must agree in only one way. Consider the following:
@@
Curiously, this definition has received relatively little attention in the hole argument literature until just recently. De re* determinism is called ‘Definition 3’ in (Belot [1995]) and ‘Dm2++’ in (Cudek [2023]). It is endorsed by Halvorson et al. ([unpublished]) who call it ‘D3’ determinism. The definition is also briefly considered by Halvorson and Manchak ([2025]) and Landsman ([2023]).
@@
Why might one be interested in de re* determinism? It has been argued that de re determinism is not strong enough to count all instances of indeterminism as such. Belot ([1995]) introduces his own symmetric example which leads to the formulation of his Definition 3 (de re* determinism). We will return to a similar example in due course. For now let us turn to another case: classical Leibnizian spacetime (Earman [1989]). One can verify that the spacetime (considered as a singleton collection) counts as deterministic under the classical analogue of the de re formulation. And yet there are reasons to think that such a spacetime ought to be considered indeterministic. Here is Earman ([1977], p. 96):
@@
If the history of a particle is represented by a timelike world line (i.e. a world line which is everywhere oblique to the planes of simultaneity) on [Leibnizian spacetime], then determinism cannot hold. For among the automorphisms [of Leibnizian space-time] are those which are the identity on the portion of [Leibnizian spacetime] on or below some given time slice but which differ from the identity above; such a mapping leaves fixed the entire past history of the particles while changing their future behavior. Since the automorphisms of the space-time should be symmetries of the dynamical laws (whatever they are), there will be two solutions which describe the same particle histories for all past times but which describe different future behaviors.
@@
Whether or not Leibnizian spacetime (or theories of matter set therein) ought to count as deterministic in some sense is an interesting question (see Stein [1977]; Weatherall [2020]). We do not enter the debate here. At this stage, we emphasize that we are not interested in capturing, once and for all, a single notion of ‘determinism’ within spacetime theories. There are many interesting forms of determinism to explore (Doboszewski [2019]; Smeenk and Wúthrich [2021]). In a pluralistic way, we are simply interested in mapping out which collections of general relativistic spacetimes are deterministic relative to which definitions. Here, we draw attention to the fact that there are definitions of determinism – in particular de re* – which are strong enough to capture Earman’s sense in which Leibnizian spacetime is indeterministic.
@@
Let us now return to general relativity. The isometry uniqueness clause which is added to de re determinism to obtain de re* determinism has the effect of requiring a certain type of spacetime asymmetry known as ‘rigidity’ (Geroch [1969]) which has recently become a focus of attention in the hole argument literature (Menon and Read [2023]; Dewar [2025]; Halvorson and Manchak [2025]). Consider the following definition:
@@
Definition 4: A collection $\mathscr{C}\subseteq\mathscr{U}$ is rigid if, for any $(M,g),(M^{\prime},g^{\prime})\in\mathscr{C}$ and any open set $O\subset M$, if there are isometries $\varphi,\psi:M\rightarrow M^{\prime}$ such that $\varphi_{|O}=\psi_{|O}$, then $\varphi=\psi$.
@@
If a collection of spacetimes is rigid, then for any pair of spacetimes in the collection, if there are isometries from one to the other that agree on an open set, they must agree entirely. It is easily verified that any collection $\mathscr{C}\subseteq\mathscr{H}$ that is both rigid and de re deterministic must be de re* deterministic. Indeed, one can think of the rigidity condition as a type of ‘dual’ to de re determinism: the latter requires the existence of a certain isometry $\psi:M\rightarrow M^{\prime}$ while the former requires (because initial segments are open sets) the uniqueness of such an isometry (Dewar [2025]). One can use a general rigidity result due to Geroch ([1969], Theorem A1) to show the following (Halvorson and Manchak [2025], Theorem 1):
Proposition 3: Any collection $\mathscr{C}\subseteq\mathscr{U}$ is rigid.
@@
One way to understand the significance of the proposition is as follows: within the context of general relativity, there can be no collection of spacetimes that is de re deterministic but not de re* deterministic (Cudek [2023], footnote 46). To distinguish between these two notions of determinism, one has to look to certain classical spacetime theories (e.g. the Leibnizian case mentioned above) or non-standard general relativity such as its non-Hausdorff generalization (Manchak and Barrett [forthcoming]). From Proposition 3, we have the following:
Corollary 1: A collection $\mathscr{C}\subseteq\mathscr{H}$ is de re deterministic if and only if it is de re* deterministic.
@@
Recall that $\mathscr{V}^{+}\subset\mathscr{H}$ is the collection of four-dimensional, inextendible, globally hyperbolic, vacuum solutions of Einstein’s equation. From Corollary 1 and Proposition 2, we have the following result which shows that significant swathes of general relativistic spacetimes of physical interest satisfy a much stronger form of determinism than previously known (Landsman [2022], footnote 23):
Corollary 2: Any collection $\mathscr{C}\subseteq\mathscr{V}^{+}$ is de re* deterministic.
We close by highlighting how de re* determinism is related to a point emphasized by Weatherall ([2018], p. 336) in his paper on the hole argument. There, a formal proposition within the context of (standard) general relativity is central: if $(M,g)$ is a spacetime and $\varphi:M\rightarrow M$ is a non-trivial hole diffeomorphism, then the identity map on $M$ is not an isometry from $(M,g)$ to $(M,\varphi_{*}g)$. Let us articulate a general version of this statement with respect to any spacetime theory $\mathscr{T}$ for which one can make sense of initial segments. Such a theory will be a collection of models of the form $(M,O_{1},...,O_{n})$ where $M$ is a manifold and $O_{1},...,O_{n}$ are geometric objects on $M$. Consider the following condition on $\mathscr{T}$.
(Id $\neq$ Iso): Let $(M,O_{1},...,O_{n})$ be model in $\mathscr{T}$ and let $U\subset M$ be an initial segment. If $\varphi:M\rightarrow M$ is a non-trivial hole diffeomorphism such that $\varphi_{|U}$ is the identity map on $U$, then the identity map on $M$ is not an isomorphism from $(M,O_{1},...,O_{n})$ to $(M,\varphi_{*}O_{1},...,\varphi_{*}O_{n})$.
How is (Id $\neq$ Iso) related to de re* determinism? If the latter is generalized in the natural way to apply to any spacetime theory $\mathscr{T}$ with initial segments, then a simple result follows: if $\mathscr{T}$ satisfies de re* determinism, then $\mathscr{T}$ must also satisfy (Id $\neq$ Iso). To see this, suppose (Id $\neq$ Iso) fails for some spacetime theory $\mathscr{T}$. So there is a model $(M,O_{1},...,O_{n})$ in $\mathscr{T}$, an initial segment $U\subset M$, and a non-trivial hole diffeomorphism $\varphi:M\rightarrow M$ such that $\varphi_{|U}$ is the identity map on $U$ and yet the identity map $\psi$ on $M$ is an isomorphism from $(M,O_{1},...,O_{n})$ to $(M,\varphi_{*}O_{1},...,\varphi_{*}O_{n})$. We know that $\varphi\neq\psi$ since $\varphi$ is non-trivial. Since the distinct maps $\varphi$ and $\psi$ are both isomorphisms from $(M,O_{1},...,O_{n})$ to $(M,\varphi_{*}O_{1},...,\varphi_{*}O_{n})$ and $\varphi_{|U}=\psi{|U}$, it follows that $\mathscr{T}$ fails the uniqueness clause in the de re* form of determinism.
Note that a similar result does not hold for (a generalized form of) de re determinism. Because this definition lacks the uniqueness clause, it is not strong enough to ensure that (Id $\neq$ Iso) is satisfied. To see this, just consider Leibnizian spacetime (taken as a singleton collection). As mentioned above, this spacetime satisfies the classical analogue of de re determinism. But since Leibnizian spacetime is not rigid, one can easily verify that it fails to satisfy (Id $\neq$ Iso).
4. De Dicto + Uniqueness